OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (3x - 1) units and the length of side ON is (x + 7) units. What value of x will make it a rhombus? 3 4 6 8

OpenStudy (anonymous):

the sides would then have to be equal 3x-1 = x+7 solve to get x = 4 test in both sides and it works (11=11) m

OpenStudy (anonymous):

Thank you! :)